OpenStudy (anonymous):

Solve the equation: -7/3n = -42 A. n=18 B. n=298 C.n=92 D. n= -21 4/7

OpenStudy (anonymous):

multiply both sides by 3n multiplying just one side by 3n/3n does nothing

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it \[\frac{ 7 }{ 3n } or \frac{ 7 }{ 3 } n\]?

OpenStudy (anonymous):

Oh, sorry. I made a mistake. You should times 3n on both sides. Then you will get -7=-42*3n. Finally -7=-126n n=7/126

OpenStudy (anonymous):